# Float Self-Tagging

### v2

@@ Abstract.
Dynamic and polymorphic languages must attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. It makes it difficult to efficiently implement IEEE754 floating-point numbers as this format does not leave an easily accessible space to store the type information. The two main floating-point number encodings in-use to this day, tagged pointers and NaN-boxing, have drawbacks. Tagged pointers entail a heap-allocation of all float objects, and NaN-boxing puts additional run time costs on type checks and the handling of all other objects.
This paper presents self-tagging, a new approach to object tagging that can attach easily accessible type information to N-bit objects while retaining the ability to use all of their N bits for data. At its core, self-tagging exploits the fact that some bit sequences appear with high probability. Superimposing tags with these frequent sequences allows encoding both N-bit data and type within a single N-bit machine word. The main application of this approach is to represent IEEE754 64-bit and 32-bit floating point numbers on 64-bit and 32-bit machines respectively. We have implemented related variants of self-tagging in one JavaScript compiler and two distinct Scheme compilers to analyze their performance and compare them to tagged pointers and NaN-boxing. Our experiments demonstrate that in practice the approach eliminates heap-allocation of IEEE754 floating-point numbers and improves the execution time of float-intensive benchmarks in Scheme by 2.4$\times$, and in JavaScript by 3.6$\times$, with a negligible performance impact on other benchmarks, which makes it a good alternative to both tagged pointers and NaN-boxing.
Dynamic Languages, Polymorphic Languages, Floating-Point, NaN-Boxing, Compiler Optimization, JavaScript, Scheme
@@ 1. Introduction
In dynamic and other polymorphic languages, efficiently handling floating-point numbers (floats) remains an elusive problem in search of an optimal solution. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses a full 64-bit word to represent a double-precision float, leaving no space for additional type information. Various representations have been developed to mitigate this issue, each with its own trade offs.
This paper presents a novel technique for encoding floats that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages. This technique does not require any program static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but also has applications on 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade offs.
@@ 1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). An N-bit machine word is used to encode an object reference (either a value or a pointer to a heap-allocated value) and a small bit field is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap-allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
@@
  • Generic pointers represent all remaining object types as pointers with a generic tag. Since the generic tag is associated with multiple types, type information is stored in the heap, typically in a header.
@@
Figure 1. The three representations in a tagged object system (here shown on a little-endian machine)
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows addition/subtraction directly on the tagged values. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. When larger integers are required, generic pointers to heap-allocated big integers can be used (Leijen, 2022). Consequently this provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
@@ 1.2. NaN-boxing
A subset of NaN’s can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaN’s, illustrated in Figure 2(b), which correspond to the interval from 0xfff8_0000_0000_0000 to 0xffff_ffff_ffff_ffff. This partitions floating-point numbers in two intervals. Bit patterns above 0xfff8_0000_0000_0000 are reserved for non-float objects. The bit patterns below 0xfff8_0000_0000_0000 are reserved for floats (including signaling NaN but not quiet NaN) and negative quiet NaN are mapped to the bit pattern 0xfff8_0000_0000_0000.
@@
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-boxing are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of all other, non-float objects due to higher-cost machine instructions necessary to check the type and dereference NaN-boxed pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. Moreover, NaN-boxing on 32-bit architectures would severely cripple memory management because 32-bit floats have NaNs that leave space for only 22-bit pointers, which only allows addressing 4 MiB.
@@ 1.3. Pointer-Biased NaN-boxing
NuN-boxing is a variant of NaN-boxing that alleviates the cost of dereferencing NaN-boxed pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaN’s greater than or equal to 0xfffe_0000_0000_0000.
@@ 1.4. Contribution
The contribution of this paper is a thorough analysis of a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects while retaining the ability to use all of their 64 bits for storing data. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrary to NaN-boxing, the newly obtained representation does not impact the performance of encoding and decoding other tagged objects. It is also applicable in the context of single-precision floats on 32-bit architectures, which NaN-boxing does not support. Self-tagging cannot represent all floats but the conducted experiments show that the set of numbers it can represent is large enough to cover all floats actually used by all surveyed JavaScript and Scheme test programs. This paper shows that with float self-tagging there are very few if any heap-allocated floats in practice, thus providing an interesting trade off between the advantages of tagging and NaN-boxing.
@@ 1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 discusses a basic form of self-tagging. Section 4 shows how to refine the approach to greatly reduce heap-allocation of floats. Section 5 presents technical details of its implementation. It shows portable implementations that could be used by compilers that generate either C or assembly code. Section 6 shows experimental results both in the context of Scheme with the Bigloo and Gambit compilers, and in the Hopc ahead-of-time JavaScript compiler. Section 7 discusses the adaptation of self-tagging to 32-bit architectures. Finally, section 8 presents related work.
@@ 2. Self-Tagging
This section describes the basic form of self-tagging, a tagging technique that exploits the fact that some values naturally contain the appropriate tag corresponding to their type, at the correct location, in their bit arrangement. While this may appear coincidental, it occurs frequently enough to allow the transformation of a significant proportion of heap allocated objects (either represented as tagged pointers or generic pointers) to tagged values. In this paper, unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging pertains to 32-bit architectures as well, which is discussed in Section 7.
Consider the common case of tagged objects with 3-bit tags and, for the sake of the example, assume that the tag 001 is chosen for floats. This setup was previously illustrated in Figure 1 where a tagged pointer refers to the heap-allocated float 3.02. The IEEE754 encoding of 3.02 happens to end with 001, the tag for floats in this example. This purely coincidental fact is bound to happen with $1/8$ of all the bit patterns in the floating-point number space, since eight distinct tags exist on 3 bits.
Self-tagging exploits occurrences where the tag bits of a pointer appear in its value. Such objects can be unboxed, making them tagged values instead of heap allocated values. However, since only $1/8$ of all floats can be unboxed in such a way, a second tag must be reserved for the remaining floats, which still need to be heap allocated with either tagged or generic pointers.
Self-tagging introduces the notion of tags that indicate which values are unboxed. This contrasts with standard object tagging that invites assigning tags with efficient handling of frequent types in mind, regardless of values. With self-tagging, the initial choice of 001 for floats ought to be reevaluated. A better choice could be 000, which captures, among other values, all integer valued floats below $12^{50}+$ 1 (the first integer valued float whose IEEE754 representation does not end in 000). The tag 001 can be kept for tagged pointer floats.
Figure 4. Representation of a (unboxed) float as a tagged value with tag 000. All floats whose IEEE754 representation ends with 000 are unboxed. Other floats are encoded as tagged pointers as usual.
Figure 4 shows the encoding of floats whose IEEE754 representation ends with the better suited tag 000 for unboxed floats. In those cases, the last three bits of the mantissa and the tag are identical. This allows superimposing them to represent both a 64-bit value and a 3-bit tag within a single machine word. Any object tagged with 000 is known to be a tagged value float. Obtaining its value requires no further transformation.
This specific example of self-tagging allows unboxing $1/8$ of floats. Yet, this proportion can be increased by assigning extra tags to unboxed floats. In general, if $T$ tags are available and $f$ tags are reserved for self-tagging of floats, then $f/T$ of all floats can be unboxed. A higher number of reserved tags $f$ can potentially lead to less memory allocation, and thus less strain on the garbage collector. However, the short supply of tags suggests that $f$ should be limited and only worthwhile tags should be selected. Tag selection is discussed in Section 4.
3. A First Experiment
This section presents the first experiment of float self-tagging with the Bigloo Scheme compiler (Manuel Serrano, 2024). Bigloo uses tagged objects by default and can be configured to use NaN-boxing and NuN-boxing (discussed in Section 6.3), making it ideal for comparing these techniques and self-tagging.
3.1. Overview of the Bigloo Compiler
The Bigloo compiler builds a typed abstract syntax tree from the compiled Scheme program. Several analysis and optimization refine the initial optional annotations contained in the source program, the main ones being occurrence-typing (Tobin-Hochstadt and Felleisen, 2010) and storage use analysis (Serrano and Feeley, 1996). During compilation, primitive values have a dual representation: one for polymorphic contexts and one for specific contexts. For instance, when it can establish that a function is always invoked with floats, the compiler assigns the C double type to that function and, when needed, it introduces cast operations from and to the generic representation of floats and double.
The result of this compilation technique is that most temporary and local functions are precisely typed and avoid polymorphic representations, which contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into vectors of C doubles. The experiment conducted in this section evaluates the impact of removing the memory allocations of floats that the compiler is not able to remove statically. For a compiler that does not implement such static analyses, the impact might be even more significant than measured with Bigloo.
3.2. Experimental Setup
Two versions of the Bigloo Scheme compiler are compared. The first is the original Bigloo version, which implements standard tagged objects where floats are heap-allocated using a tagged pointer. The second implementation only differs by its use of self-tagged floats.
In the self-tagging variant, the two tags 000 and 100 are reserved for self-tagged floats, thus increasing the proportion of self-tagged floats to $1/4$. This means that all floats $n$ that are integers such that $|n|\leq 2^{51}$ as well as those values scaled by a power of two, will be unboxed floats. The tag 001 is reserved for the remaining floats, which are represented by tagged pointers.
Both implementations are tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between benchmarks that use few or no floats and float-intensive benchmarks (fibfp=doubly recursive Fibonacci with floats, fft=1024 point Fast Fourier Transform, mbrot=mandelbrot set, nucleic=3D structure determination of a nucleic acid, pnpoly=determine if a 2D point is in a 20 sided polygon, ray=raytrace a scene with 33 spheres, simplex=linear programming using simplex method, sum1=add a file of floats, and sumfp=add the floats 0 to $10^{6}$).
The following results were obtained on a machine with an Intel Core i7-7700K, 48 GB of RAM, and under Debian 10.13 with kernel version SMP Debian 4.19.269-1.
3.3. Memory Allocation
(a) Relative memory allocation
(b) Relative execution time
(c) Relative missed branch prediction
Figure 5. Heap allocations, relative execution time and missed branch predictions comparison between the original Bigloo version and the self-tagging version. The y-axis shows each measure relative to the original version. 1.0 indicates no change. For all metrics, lower is better.
Memory usage is measured by counting the number of heap-allocated bytes across the lifetime of each benchmark. The original and self-tagging Bigloo versions are then compared to assess the efficiency of self-tagging at unboxing floats and its impact on benchmarks that make limited use of floats. Results are shown in Figure 5(a).
On all benchmarks extensively using floats, memory allocation decreases. On mbrot, nucleic, simplex and sum1, allocations decrease by about 25%, which is consistent with the proportion of self-tagged floats ($1/4$). The fft benchmark has only 17% fewer allocations due to predominantly using the float 0.0, which the original Bigloo version preallocates. For fibfp, pnpoly, ray and sumfp there is very little allocation of floats when using self-tagging because those programs mostly use “simple” floats that are integer valued floats $\leq 2^{51}$ scaled by a small power of two, such as 0.75, 2.25, or 42.0.
On all benchmarks that use few floats, memory usage increases by less than 1%. This tiny increase stems from the use of generic pointers for types that were assigned specific tags in the original version.
3.4. Execution Time
Figure 5(b) shows the execution time comparison between both implementations. Execution time refers to elapsed real time measured with perf-stat. Each benchmark is parameterized for a run to last for at least 10 seconds, and is repeated 5 times. All timing results are the average execution time of these runs.
On a geometric average, float-intensive benchmarks are about 25% faster while non-float benchmarks are unaffected. However, the relation between lower memory usage and faster execution is noisy. The execution of mbrot and nucleic is significantly slower despite allocating about $1/4$ less floats with self-tagging.
This unpredictability is explainable, and fixable. Figure 5(c) shows that, despite reducing float allocations in mbrot and nucleic, self-tagging also causes about 4 times more missed branch predictions. This stems from the fact that tags occupy the low bits of the mantissa. In that position, the smallest variations cause a float’s representation to switch between tagged value and tagged pointer in a hard to predict way. This is beyond the capabilites of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
@@ 3.4. Execution Time
bits bits + 1 range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1
0.0 94% 13% 0% 0% 16% 13% 11% 0% 0%
00000 00001 5e-324 .. 2.1e-289 - - - - - - - - -
00001 00010 2.1e-289 .. 3.8e-270 - - - - - - - - -
00010 00011 3.8e-270 .. 7e-251 - - - - - - - - -
00011 00100 7e-251 .. 1.3e-231 - - - - - - - - -
00100 00101 1.3e-231 .. 2.4e-212 - - - - - - - - -
00101 00110 2.4e-212 .. 4.4e-193 - - - - - - - - -
00110 00111 4.4e-193 .. 8.1e-174 - - - - - - - - -
00111 01000 8.1e-174 .. 1.5e-154 - - - - - - - - -
01000 01001 1.5e-154 .. 2.8e-135 - - - - - - - - -
01001 01010 2.8e-135 .. 5.1e-116 - - - - - - - - -
01010 01011 5.1e-116 .. 9.4e-97 - - - - - - - - -
01011 01100 9.4e-97 .. 1.7e-77 - - - - - - - - -
01100 01101 1.7e-77 .. 3.2e-58 - - - - - - - - -
01101 01110 3.2e-58 .. 5.9e-39 - - - - - - - - -
01110 01111 5.9e-39 .. 1.1e-19 - - - - - - - - -
01111 10000 1.1e-19 .. 2 6% 33% 96% 70% 60% 18% 62% 0% 0%
10000 10001 2 .. 3.7e19 0% 54% 4% 30% 24% 69% 27% 100% 100%
10001 10010 3.7e19 .. 6.8e38 - - - - - - - - -
10010 10011 6.8e38 .. 1.3e58 - - - - - - - - -
10011 10100 1.3e58 .. 2.3e77 - - - - - - - - -
10100 10101 2.3e77 .. 4.3e96 - - - - - - - - -
10101 10110 4.3e96 .. 7.9e115 - - - - - - - - -
10110 10111 7.9e115 .. 1.5e135 - - - - - - - - -
10111 11000 1.5e135 .. 2.7e154 - - - - - - - - -
11000 11001 2.7e154 .. 4.9e173 - - - - - - - - -
11001 11010 4.9e173 .. 9.1e192 - - - - - - - - -
11010 11011 9.1e192 .. 1.7e212 - - - - - - - - -
11011 11100 1.7e212 .. 3.1e231 - - - - - - - - -
11100 11101 3.1e231 .. 5.7e250 - - - - - - - - -
11101 11110 5.7e250 .. 1.1e270 - - - - - - - - -
11110 11111 1.1e270 .. 1.9e289 - - - - - - - - -
1.9e289 .. 1.8e308 - - - - - - - - -
11111 00000 Infinity/NaN - - - - - - - - -
@@
Figure 6. The positive value ranges captured by each combination of the 5 most significant exponent bits and the proportion of all floats computed by the float-intensive programs in the R7RS benchmark suite. Intervals of interest are highlighted with a background color. Empty entries indicate no float in this range was computed, whereas 0% means that very few floats in this range were generated (less than 0.5%).
4. Improved Predictability and Coverage
This section addresses the issue uncovered by the previous experiment and presents an improved, performant variant of float self-tagging.
The previous experiment shows that while self-tagging decreases the number of heap-allocated floats it does not consistently improve performance due to failed branch predictions. To address this issue, self-tagging is used for the most common floats encountered in practice to maximize the likelihood that a float will be self-tagged, thus improving branch prediction and decreasing heap-allocated floats.

### v3

@@ Abstract.
Dynamic and polymorphic languages attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. This makes it difficult to efficiently implement IEEE754 floating-point numbers as this format does not leave an easily accessible space to store type information. The three main floating-point number encodings in use today, tagged pointers, NaN-boxing, and NuN-boxing, have drawbacks. Tagged pointers entail a heap allocation of all float objects, and NaN/NuN-boxing puts additional run time costs on type checks and the handling of other objects.
This paper introduces self-tagging, a new approach to object tagging that uses an invertible bitwise transformation to map floating-point numbers to tagged values that contain the correct type information at the correct position in their bit pattern, superimposing both their value and type information in a single machine word. Such a transformation can only map a subset of all floats to correctly typed tagged values, hence self-tagging takes advantage of the non-uniform distribution of floating point numbers used in practice to avoid heap allocation of the most frequently encountered floats.
Variants of self-tagging were implemented in two distinct Scheme compilers and evaluated on four microarchitectures to assess their performance and compare them to tagged pointers, NaN-boxing, and NuN-boxing. Experiments demonstrate that, in practice, the approach eliminates heap allocation of nearly all floating-point numbers and provides good execution speed of float-intensive benchmarks in Scheme with a negligible performance impact on other benchmarks, making it an attractive alternative to tagged pointers, alongside NaN-boxing and NuN-boxing.
Dynamic Languages, Polymorphic Languages, Floating-Point Numbers, NaN-Boxing, Compiler Optimization, Scheme
@@ 1. Introduction
In dynamic and other polymorphic languages, efficiently supporting floating-point numbers (floats) remains a challenging issue. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses the entire 64-bit word to represent a double-precision float, leaving no space for additional type information. A straightforward but suboptimal solution is to represent floats as heap allocated objects. To reduce the overhead of heap allocating floats, several alternative representations have been proposed, each offering different trade-offs.
This paper presents a novel technique for encoding floats, called self-tagging, that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages.
Self-tagging exploits the fact that any invertible bitwise transformation on an IEEE754 float will map some floats to a bit arrangement that contains the correct type information at the correct location, thus requiring no memory allocation for these floats. Using the fact that some ranges of floats appear more frequently in practical applications (Lindstrom and Isenburg, 2006; Brown et al., 2007; Afroozeh et al., 2023; Zurstrassen et al., 2023), it is possible to define a transformation that avoids heap allocating almost all floats encountered in practice.
This technique does not require any program static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but is also applicable to 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade-offs.
@@ 1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). An N-bit machine word is used to encode an object reference (either a value or a pointer to a heap allocated value) and a small bit field is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
@@
  • Generic pointers represent all remaining objects as pointers with a generic tag. The generic tag is associated with multiple types, and type information is stored in a header, in the heap.
@@
Figure 1. The three representations in a tagged object system (here shown on a little-endian machine).
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows direct addition/subtraction. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade-off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. For larger values, generic pointers to heap allocated big integers can be used (Leijen, 2022). This provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
@@ 1.2. NaN-boxing
A subset of NaNs can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaNs, illustrated in Figure 2(b), which correspond to the interval from 0xfff8_0000_0000_0000 to 0xffff_ffff_ffff_ffff. This partitions floating-point numbers in two intervals. Bit patterns above 0xfff8_0000_0000_0000 are reserved for non-float objects. The bit patterns below 0xfff8_0000_0000_0000 are reserved for floats (including signaling NaN but not quiet NaN) and negative quiet NaN are mapped to the bit pattern 0xfff8_0000_0000_0000.
@@
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-boxing are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of other, non-float objects due to higher-cost machine instructions to check the type and dereference NaN-boxed pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. Moreover, on 32-bit architectures, NaN-boxing would cripple memory management because 32-bit float NaNs leave space for only 22-bit pointers, which only allows addressing 4 MiB.
@@ 1.3. Pointer-Biased NaN-boxing
NuN-boxing is a variant of NaN-boxing that alleviates the cost of dereferencing NaN-boxed pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaNs greater than or equal to 0xfffe_0000_0000_0000.
@@ 1.4. Contribution
The contribution of this paper is a thorough analysis of a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects by applying a bitwise transformation that preserves their value without heap allocation. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrarily to NaN-boxing, it does not impact the performance of encoding and decoding other tagged objects and is applicable in the context of 32-bit architectures. Self-tagging cannot represent all floats and thus requires an hybrid representation for fast self-tagged and slow heap allocated floats. Yet, the conducted experiments show that the set of numbers covered by self-tagging is large enough to remove nearly all heap allocations of floats in practice. Hence, it provides another alternative to allocated floats with different trade-offs than NaN/NuN-boxing.
@@ 1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 presents technical details for portable implementations that could be used by compilers that generate either C or assembly code. Section 4 shows experimental results in the context of Scheme with the Bigloo and Gambit compilers. Section 5 discusses the adaptation of self-tagging to 32-bit architectures. Finally, Section 6 presents related work.
@@ 2. Self-Tagging
This section describes self-tagging, a tagging technique that exploits the fact that some values can be encoded to a tagged value by a bitwise transformation containing the tag corresponding to their type. Such self-tagged objects avoid the cost of heap allocation. In this paper, self-tagging is applied to the specific case of IEEE754 floats. Unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging also pertains to 32-bit architectures, which is discussed in Section 5.
Consider the common case of tagged objects with 3-bit tags (8 available tags). Given tag $T$ for floats, any invertible bitwise transformation will map 1/81/81 / 8 of possible floats to a bit arrangement that contains $T$ on its low bits. These floats can thus be encoded to tagged values with this transformation and decoded back to a IEEE754 representation by applying the inverse transformation. The set of floats that are mapped to a correctly tagged value by a given transformation are said to be self-tagged and require no memory allocation. Since only a subset of all floats can be self-tagged, the remaining floats must be heap allocated with tagged or generic pointers. Yet, the proportion of self-tagged floats can be increased by assigning more than one tag to self-tagged floats. In general, using $n$ tags for self-tagging allows avoiding heap allocation of $n/8$ of all possible floats.
As an example, consider the case where the encoding transformation is a 4-bit rotation to the left, which has the effect of using the 3 high bits of an IEEE754 float’s exponent as tag for self-tagged floats (Figure 4). This transformation has the interesting property that each tag captures a contiguous range of floats. For instance, assigning the tags 000, 011 and 100 to self-tagged floats, avoids heap allocating all floats in the ranges $0$. $7\times 10^{-251}$ and $1.7\times 10^{-77}$. $2.3\times 10^{77}$ as well as the corresponding negative ranges (detailed calculations of these ranges are provided in Section 2.1).
This suggests that, while many transformations can be chosen to implement self-tagging, some choices are better than others. Three properties make for a useful self-tagging scheme. It must:
  • Capture intervals of common floats: most values computed in practice are either ±0.0$\pm$ 0.0± 0.0 or centered around ±1.0$\pm$ 1.0± 1.0 (Lindstrom and Isenburg, 2006; Brown et al., 2007; Afroozeh et al., 2023; Zurstrassen et al., 2023). These should be prioritized to reduce memory allocation.
  • Have an efficient encoding/decoding: the cost of the transformation must not outweigh the benefit of decreased memory allocations and strain on the garbage collector.
  • Use few tags: it should reserve as few tags as possible to avoid having to defer to generic pointers for other important types. Ideally, a single tag should be used for all common floats, and uncommon floats should be represented either with tagged or generic pointers.
Section 2.1 presents self-tagging variants that offer different tradeoffs between captured intervals and number of reserved tags. Section 3 describes efficient implementations for each variant.
@@
bits bits + 1 range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1 (Zurstrassen et al., 2023)
0.0 79% 9% 0% 1% 11% 7% 11% 13% 0% 0%
00000 00001 5e-324 .. 2.1e-289 - - - - - - - - - 0%
00001 00010 2.1e-289 .. 3.8e-270 - - - - - - - - - 0%
00010 00011 3.8e-270 .. 7e-251 - - - - - - - - - 0%
00011 00100 7e-251 .. 1.3e-231 - - - - - - - - - 0%
00100 00101 1.3e-231 .. 2.4e-212 - - - - - - - - - 0%
00101 00110 2.4e-212 .. 4.4e-193 - - - - - - - - - 0%
00110 00111 4.4e-193 .. 8.1e-174 - - - - - - - - - 0%
00111 01000 8.1e-174 .. 1.5e-154 - - - - - - - - - 0%
01000 01001 1.5e-154 .. 2.8e-135 - - - - - - - - - 0%
01001 01010 2.8e-135 .. 5.1e-116 - - - - - - - - - 0%
01010 01011 5.1e-116 .. 9.4e-97 - - - - - - - - - 0%
01011 01100 9.4e-97 .. 1.7e-77 - - - - - - - - - 0%
01100 01101 1.7e-77 .. 3.2e-58 - - - - - - - - - 0%
01101 01110 3.2e-58 .. 5.9e-39 - - - - - - - - - 1%
01110 01111 5.9e-39 .. 1.1e-19 0% - - - - 0% - - - 3%
01111 10000 1.1e-19 .. 2 21% 28% 89% 61% 63% 22% 64% 13% 0% 61%
10000 10001 2 .. 3.7e19 0% 63% 11% 38% 26% 70% 25% 75% 100% 28%
10001 10010 3.7e19 .. 6.8e38 - - - - - - - - - 2%
10010 10011 6.8e38 .. 1.3e58 - - - - - - - - - 1%
10011 10100 1.3e58 .. 2.3e77 - - - - - - - - - 0%
10100 10101 2.3e77 .. 4.3e96 - - - - - - - - - 0%
10101 10110 4.3e96 .. 7.9e115 - - - - - - - - - 0%
10110 10111 7.9e115 .. 1.5e135 - - - - - - - - - 0%
10111 11000 1.5e135 .. 2.7e154 - - - - - - - - - 0%
11000 11001 2.7e154 .. 4.9e173 - - - - - - - - - 0%
11001 11010 4.9e173 .. 9.1e192 - - - - - - - - - 0%
11010 11011 9.1e192 .. 1.7e212 - - - - - - - - - 0%
11011 11100 1.7e212 .. 3.1e231 - - - - - - - - - 0%
11100 11101 3.1e231 .. 5.7e250 - - - - - - - - - 0%
11101 11110 5.7e250 .. 1.1e270 - - - - - - - - - 0%
11110 11111 1.1e270 .. 1.9e289 - - - - - - - - - 0%
1.9e289 .. 1.8e308 - - - - - 0% - - - 0%
11111 00000 Infinity/NaN - - - - - - - - - N/A
@@
Figure 5. The absolute value ranges captured by each combination of the 5 most significant exponent bits and the proportion of all floats computed by float-intensive R7RS benchmarks. Intervals of interest are highlighted. Empty entries indicate no float in this range, whereas 0% means that very few floats were generated (less than 0.5%). The last column shows the relative occurence of float for each interval reported by (Zurstrassen et al., 2023, Fig. 7).
2.1. Float Coverage
Self-tagging introduces the notion of tags that indicate which values of a type are not heap allocated. This contrasts with standard object tagging that invites assigning tags with efficient handling of frequent types in mind, regardless of values. Consequently, self-tagging should be used for the most common floats to maximize the likelihood that a float will be self-tagged. Floats whose magnitude is concentrated near 1.01.01.0 have been observed to be more common (Brown et al., 2007; Lindstrom and Isenburg, 2006; Afroozeh et al., 2023; Zurstrassen et al., 2023), with a relative occurrence that decreases sharply away from 1.01.01.0 (Zurstrassen et al., 2023). Zeros are common, but subnormal numbers are rarely used (Zurstrassen et al., 2023).
These observations from (Brown et al., 2007; Lindstrom and Isenburg, 2006; Afroozeh et al., 2023; Zurstrassen et al., 2023) are reproduced by instrumenting the R7RS benchmarks, the standard benchmark suite for Scheme (r7r, 2024), to create a profile of computed floats. All benchmarks labelled as float benchmarks were profiled (this includes fibfp=recursive Fibonacci with floats, fft=1024 points Fast Fourier Transform, mbrot=mandelbrot set, nucleic=3D structure determination of a nucleic acid, pnpoly=determine if a 2D point is in a 20 sided polygon, ray=raytrace a scene of 33 spheres, simplex=linear programming using simplex method, sumfp=add the floats 0 to $10^{6}$, and sum1=add a file of floats).